# Carbon and Silicon, Coexist or Compete? A Survey on Human-AI Interactions in Agent-based Modeling and Simulation

### v1

@@ 5.2. Goal 2: Explore Different Scenarios
2023 | Generative Agents (Park et al., 2023b) | S-P | D-S | Actor | Agent ![]() |
@@
2022 | Social Simulacra (Park et al., 2022) | S-V | D-S | Director | Sim ![]() |
@@
2023 | ChatEval (Chan et al., 2023) | None | Pre-S | Director | Env ![]() |
@@
2023 | AgentSims (Lin et al., 2023) | S-P | D-S | Actor | Agent ![]() |
@@
2024 | AgentCoord (Pan et al., 2024) | S-P | D-S | Director | Agent ![]() ![]() |
@@
2024 | Rehearsal (Shaikh et al., 2024) | None | D-S | Director | Agent ![]() |
@@
2023 | Humanoid Agents (Wang et al., 2023a) | S-P | D-S | Director | Sim ![]() |
@@
2024 | Zhang et al. (Zhang et al., 2024d) | R-V | D-S | Director | Agent ![]() |
@@
2024 | ChatCam (Xiao et al., 2024) | R-P | D-S | Director | Agent ![]() |
@@
2024 | Cuadra et al. (Cuadra et al., 2024) | R-P | D-S | Director | Agent ![]() |
@@
2024 | CrowdBot (Wang et al., 2024b) | R-P | D-S | Director | Agent ![]() |
@@
2024 | Sheshadri et al. (Sheshadri and Hara, 2024) | R-P | D-S | Director | Agent ![]() |
@@ 5.5. Goal 5: Analyze Simulation Data
- | AnyLogic (Borshchev, 2014) | S-P | D-S | Observer | Agent ![]() |
@@
- | EINSTein (Berryman, 2008) | S-P | Post-S | Observer | Agent ![]() |
@@
2006 | MANA (Berryman, 2008) | S-P | Post-S | Observer | Agent ![]() |
@@
1999 | NetLogo (Wilensky, 1999) | S-P | Post-S | Observer | Agent ![]() |
@@
2006 | North et al. (North et al., 2006) | S-P | Post-S | Observer | Agent ![]() |
@@
1996 | Swarm (Minar et al., 1996) | S-P | D-S | Observer | Agent ![]() |
@@
2023 | Zarzà et al. (de Zarzà et al., 2023) | S-P | Post-S | Observer | Agent ![]() |
@@
2024 | AgentLens (Lu et al., 2024b) | S-P | Post-S | Observer | Agent ![]() |
@@
2024 | AgentCoord (Pan et al., 2024) | S-P | D-S | Observer | Agent ![]() ![]() |
@@
2023 | Humanoid Agents (Wang et al., 2023a) | S-P | Post-S | Observer | Agent ![]() |
@@ Appendix A Appendix
2023 | Generative Agents (Park et al., 2023b) | S-P | Pre-S | Scriptwriter |
Agent ![]() |
@@
2022 | Social Simulacra (Park et al., 2022) | S-V | Pre-S | Scriptwriter |
Agent ![]() ![]() ![]() |
@@
2023 | ChatEval (Chan et al., 2023) | None | Pre-S | Scriptwriter |
Agent ![]() ![]() |
@@
2023 | MetaGPT (Hong et al., 2024) | None | Pre-S | Scriptwriter |
Agent ![]() |
@@
2023 | Argyle et al. (Argyle et al., 2023) | R-P | Pre-S | Prototype |
Agent ![]() |
@@
2023 | SayPlan (Rana et al., 2023) | S-P | Pre-S | Director |
Agent ![]() |
@@
2023 | AgentSims (Lin et al., 2023) | S-P | Pre-S | Scriptwriter |
Agent ![]() ![]() ![]() |
@@
2022 | Huang et al. (Huang et al., 2022) | S-P; R-P | Pre-S | Director |
Agent ![]() |
@@
2022 | WebShop (Yao et al., 2022) | S-V | Pre-S | Director |
Agent ![]() |
@@
2023 | Mind2Web (Deng et al., 2023) | R-V | Pre-S | Director |
Agent ![]() |
@@
2023 | CAMEL (Li et al., 2023b) | None | Pre-S | Director |
Agent ![]() |
@@
2023 | Aher et al. (Aher et al., 2023) | R-P | Pre-S | Prototype |
Agent ![]() |
@@
2023 | CGMI (Jinxin et al., 2023) | S-P | Pre-S | Scriptwriter |
Env ![]() |
@@
2023 | ChatLaw (Cui et al., 2024b) | None | Pre-S | Director |
Agent ![]() |
@@
2020 | Alfred (Shridhar et al., 2020) | R-P | Pre-S | Director |
Agent ![]() |
@@
2023 | ChatDev (Qian et al., 2024) | None | Pre-S | Director |
Agent ![]() |
@@
- | BactoWars (Berryman, 2008) | S-P | Pre-S | Scriptwriter |
Agent ![]() ![]() ![]() |
@@
- | EINSTein (Berryman, 2008) | S-P | Pre-S | Scriptwriter |
Agent ![]() ![]() ![]() ![]() |
@@
- | MANA (Berryman, 2008) | S-P | Pre-S | Scriptwriter |
Env ![]() ![]() |
@@
2005 | MASON (Luke et al., 2005) | S-P | Pre-S | Scriptwriter |
Agent ![]() ![]() ![]() |
@@
1999 | NetLogo (Wilensky, 1999) | S-P | Pre-S | Scriptwriter |
Agent ![]() ![]() |
@@
2006 | North et al. (North et al., 2006) | S-P | Pre-S | Scriptwriter |
Agent ![]() ![]() ![]() |
@@
1996 | Minar et al. (Minar et al., 1996) | S-P | Pre-S | Scriptwriter |
Agent ![]() ![]() ![]() ![]() |
@@
2023 | ComplexWorld (Basavatia et al., 2023) | S-V | Pre-S | Scriptwriter |
Env ![]() ![]() ![]() |
@@
2013 | Gaube et al. (Gaube and Remesch, 2013) | R-P | Pre-S | Prototype |
Agent ![]() |
@@
2023 | WarAgent (Hua et al., 2024) | S-P | Pre-S | Scriptwriter |
Agent ![]() ![]() ![]() ![]() |
@@
2018 | Kavak et al. (Kavak et al., 2018) | S-P | Pre-S | Prototype |
Agent ![]() |
@@
2023 | Modelscope-agent (Li et al., 2023a) | R-V | Pre-S | Scriptwriter |
Agent ![]() ![]() |
@@
2024 | AgentCoord (Pan et al., 2024) | S-P | Pre-S | Director |
Agent ![]() |
@@
2023 | Choicemates (Park et al., 2023a) | None | Pre-S | Director |
Agent ![]() ![]() |
@@
2024 | Rehearsal (Shaikh et al., 2024) | None | Pre-S | Director |
Env ![]() |
@@
2023 | Wang et al. (Wang et al., 2024c) | S-V | Pre-S | Scriptwriter |
Agent ![]() |
@@
2023 | Humanoid Agents (Wang et al., 2023a) | S-P | Pre-S | Scriptwriter |
Agent ![]() |
@@
- | PedSim (PedSim, [n. d.]) | S-P | Pre-S | Scriptwriter |
Agent ![]() ![]() ![]() ![]() |
@@
- | AnyLogic (Borshchev, 2014) | S-P | Pre-S | Scriptwriter |
Agent ![]() ![]() ![]() ![]() |
@@
2023 | AutoGPT (Significant Gravitas, [n. d.]) | None | Pre-S | Scriptwriter |
Agent ![]() |
@@
2023 | BabyAGI (BabyAGI, [n. d.]) | None | Pre-S | Scriptwriter |
Agent ![]() |
@@
2023 | CommunityBots (Jiang et al., 2023) | None | Pre-S | Director |
Agent ![]() |
@@
2024 | ComPeer (Liu et al., 2024b) | None | Pre-S | Director |
Agent ![]() |
@@
2024 | PrISM-Observer (Arakawa et al., 2024) | R-P | Pre-S | Director |
Agent ![]() |
@@
2024 | Jaber et al. (Jaber et al., 2024) | R-P | Pre-S | Director |
Agent ![]() |
@@
2024 | Chen et al. (Chen et al., 2024) | None | Pre-S | Director |
Agent ![]() |
@@
2024 | Zhang et al. (Zhang et al., 2024d) | R-V | Pre-S | Director |
Agent ![]() |
@@
2024 | Text2AC (Sun et al., 2024) | R-V | Pre-S | Director |
Agent ![]() |
@@
2023 | Ross et al. (Ross et al., 2023) | None | Pre-S | Director |
Agent ![]() |
@@
2024 | ChatCam (Xiao et al., 2024) | R-P | Pre-S | Director |
Agent ![]() |
@@
2024 | DrHouse (Yang et al., 2024) | R-P | Pre-S | Director |
Agent ![]() |
@@
2024 | ChatIoT (Gao et al., 2024b) | R-P | Pre-S | Director |
Agent ![]() |
@@
2024 | CrowdBot (Wang et al., 2024b) | R-P | Pre-S | Director |
Agent ![]() |
@@
2022 | Social Simulacra (Park et al., 2022) | S-V | D-S | Scriptwriter | Agent ![]() ![]() |
@@
2012 | Prom Week (McCoy et al., 2012) | S-P | D-S | Director | Agent ![]() |
@@
2011 | Prom Week (McCoy et al., 2011) | S-P | D-S | Director | Agent ![]() |
@@
2023 | AGENTVERSE (Chen et al., 2023) | S-P | D-S | Director | Sim ![]() |
@@
2023 | ChatEval (Chan et al., 2023) | None | D-S | Director | Agent ![]() ![]() |
@@
2024 | Zhang et al. (Zhang et al., 2024c) | S-P | D-S | Actor | Agent ![]() |
@@
2023 | Memory sandbox (Huang et al., 2023) | None | D-S | Scriptwriter | Agent ![]() |
@@
2022 | Inner monologue (Huang et al., 2022) | S-P; R-P | D-S | Director | Agent ![]() |
@@
2022 | Krishna et al. (Krishna et al., 2022) | R-V | Pre-S | Director | Agent ![]() |
@@
2023 | RoCo (Mandi et al., 2023) | R-P | D-S | Actor | Agent ![]() |
@@
2023 | ChatLaw (Cui et al., 2024b) | None | D-S | Director | Agent ![]() |
@@
2023 | Mehta et al. (Mehta et al., 2024) | S-P | D-S | Director | Agent ![]() |
@@
2020 | Alfred (Shridhar et al., 2020) | R-P | D-S | Director | Agent ![]() |
@@
2023 | Mohanty et al. (Mohanty et al., 2023) | S-P | D-S | Director | Agent ![]() |
@@
2022 | Teach (Padmakumar et al., 2022) | R-P | D-S | Director | Agent ![]() |
@@
2005 | MASON (Luke et al., 2005) | S-P | D-S | Director | Sim ![]() |
@@
2006 | North et al. (North et al., 2006) | S-P | D-S | Director | Sim ![]() |
@@
2023 | Drive like a human (Fu et al., 2023) | R-P | D-S | Director | Agent ![]() |
@@
2006 | Guyot et al. (Guyot and Honiden, 2006) | R-V | D-S | Actor | Agent ![]() |
@@
2023 | Surrealdriver (Jin et al., 2024b) | S-P | Pre-S | Director | Agent ![]() |
@@
2023 | The SocialAI School (Kovač et al., 2023) | S-P | Pre-S | Director | Agent ![]() |
@@
2023 | Choicemates (Park et al., 2023a) | None | D-S | Director | Agent ![]() ![]() |
@@
2023 | Ghost in the minecraft (Zhu et al., 2023) | R-V | D-S | Director | Agent ![]() |
@@
2024 | Lu et al. (Lu et al., 2024a) | None | D-S | Director | Agent ![]() |
@@
2024 | Teach AI How to Code (Jin et al., 2024a) | None | D-S | Director | Agent ![]() |
@@
2024 | Zhou et al. (Zhou et al., 2024a) | None | D-S | Director | Agent ![]() |
@@
2022 | Wordcraft (Yuan et al., 2022) | None | D-S | Director | Agent ![]() |
@@
2023 | Ross et al. (Ross et al., 2023) | None | D-S | Director | Sim ![]() |
@@
2023 | Generative Agents (Park et al., 2023b) | S-P | Post-S | Actor | Agent ![]() ![]() ![]() |
@@
- | - | - | Post-S | Observer | Agent ![]() |
@@
2022 | Social Simulacra (Park et al., 2022) | S-V | Pre-S | Scriptwriter | Agent ![]() ![]() |
@@
- | - | - | Post-S | Observer | Agent ![]() |
@@
2024 | Zhang et al. (Zhang et al., 2024c) | S-P | D-S | Actor | Agent ![]() |
@@
2023 | MetaGPT (Hong et al., 2024) | None | Post-S | Observer | Agent ![]() |
@@
2023 | Argyle et al. (Argyle et al., 2023) | R-P | Post-S | Observer | Agent ![]() |
@@
2023 | AgentSims (Lin et al., 2023) | S-P | D-S | Observer | Agent ![]() |
@@
2023 | Saha et al. (Saha et al., 2023) | None | D-S | Actor | Agent ![]() |
@@
2023 | CAMEL (Li et al., 2023b) | None | Post-S | Observer | Agent ![]() |
@@
- | BactoWars (Berryman, 2008) | S-P | Post-S | Observer | Agent ![]() |
@@
2022 | FAIR et al. († et al.(2022), FAIR) | R-V | D-S | Actor | Agent ![]() |
@@
2020 | Feng et al. (Feng et al., 2020) | R-P | Pre-S | Prototype | Agent ![]() |
@@
2023 | Hämäläinen et al. (Hämäläinen et al., 2023) | R-P | Post-S | Observer | Agent ![]() |
@@
2023 | War and Peace (Hua et al., 2024) | S-P | Post-S | Observer | Agent ![]() |
@@
2023 | Surrealdriver (Jin et al., 2024b) | S-P | Post-S | Observer | Agent ![]() |
@@
2023 | Modelscope-agent (Li et al., 2023a) | R-V | Pre-S | Observer | Agent ![]() |
@@
2023 | Liu et al. (Liu et al., 2023) | S-P | Post-S | Observer | Agent ![]() |
@@
2023 | Alympics (Mao et al., 2024) | S-V | Post-S | Observer | Agent ![]() |
@@
2024 | AgentCoord (Pan et al., 2024) | S-P | D-S | Director | Agent ![]() |
@@
2023 | Choicemates (Park et al., 2023a) | None | D-S | Observer | Agent ![]() |
@@
- | - | - | Post-S | Observer | Agent ![]() |
@@
2024 | Schwitzgebel et al. (Schwitzgebel et al., 2024) | None | Post-S | Observer | Agent ![]() |
@@
2024 | Rehearsal (Shaikh et al., 2024) | None | D-S | Director | Agent ![]() |
@@
2023 | Wang et al. (Wang et al., 2024c) | S-V | Post-S | Observer | Agent ![]() |
@@
2023 | Humanoid Agents (Wang et al., 2023a) | S-P | D-S | Observer | Agent ![]() |
@@
- | - | - | Post-S | Observer | Agent ![]() |

### v2

@@ 5.2. Goal 2: Explore Different Scenarios
| 2023 | Generative Agents (Park et al., 2023b) | S-P | D-S | Actor | Agent |
@@
| 2022 | Social Simulacra (Park et al., 2022) | S-V | D-S | Director | Sim |
@@
| 2023 | ChatEval (Chan et al., 2023) | None | Pre-S | Director | Env |
@@
| 2023 | AgentSims (Lin et al., 2023) | S-P | D-S | Actor | Agent |
@@
| 2024 | AgentCoord (Pan et al., 2024) | S-P | D-S | Director | Agent |
@@
| 2024 | Rehearsal (Shaikh et al., 2024) | None | D-S | Director | Agent |
@@
| 2023 | Humanoid Agents (Wang et al., 2023a) | S-P | D-S | Director | Sim |
@@
| 2024 | Zhang et al. (Zhang et al., 2024d) | R-V | D-S | Director | Agent |
@@
| 2024 | ChatCam (Xiao et al., 2024) | R-P | D-S | Director | Agent |
@@
| 2024 | Cuadra et al. (Cuadra et al., 2024) | R-P | D-S | Director | Agent |
@@
| 2024 | CrowdBot (Wang et al., 2024b) | R-P | D-S | Director | Agent |
@@
| 2024 | Sheshadri et al. (Sheshadri and Hara, 2024) | R-P | D-S | Director | Agent |
@@ 5.5. Goal 5: Analyze Simulation Data
| - | AnyLogic (Borshchev, 2014) | S-P | D-S | Observer | Agent |
@@
| - | EINSTein (Berryman, 2008) | S-P | Post-S | Observer | Agent |
@@
| 2006 | MANA (Berryman, 2008) | S-P | Post-S | Observer | Agent |
@@
| 1999 | NetLogo (Wilensky, 1999) | S-P | Post-S | Observer | Agent |
@@
| 2006 | North et al. (North et al., 2006) | S-P | Post-S | Observer | Agent |
@@
| 1996 | Swarm (Minar et al., 1996) | S-P | D-S | Observer | Agent |
@@
| 2023 | Zarzà et al. (de Zarzà et al., 2023) | S-P | Post-S | Observer | Agent |
@@
| 2024 | AgentLens (Lu et al., 2024b) | S-P | Post-S | Observer | Agent |
@@
| 2024 | AgentCoord (Pan et al., 2024) | S-P | D-S | Observer | Agent |
@@
| 2023 | Humanoid Agents (Wang et al., 2023a) | S-P | Post-S | Observer | Agent |
@@ Appendix A Appendix
| 2023 | Generative Agents (Park et al., 2023b) | S-P | Pre-S | Scriptwriter |
Agent |
@@
| 2022 | Social Simulacra (Park et al., 2022) | S-V | Pre-S | Scriptwriter |
Agent |
@@
| 2023 | ChatEval (Chan et al., 2023) | None | Pre-S | Scriptwriter |
Agent |
@@
| 2023 | MetaGPT (Hong et al., 2024) | None | Pre-S | Scriptwriter |
Agent |
@@
| 2023 | Argyle et al. (Argyle et al., 2023) | R-P | Pre-S | Prototype |
Agent |
@@
| 2023 | SayPlan (Rana et al., 2023) | S-P | Pre-S | Director |
Agent |
@@
| 2023 | AgentSims (Lin et al., 2023) | S-P | Pre-S | Scriptwriter |
Agent |
@@
| 2022 | Huang et al. (Huang et al., 2022) | S-P; R-P | Pre-S | Director |
Agent |
@@
| 2022 | WebShop (Yao et al., 2022) | S-V | Pre-S | Director |
Agent |
@@
| 2023 | Mind2Web (Deng et al., 2023) | R-V | Pre-S | Director |
Agent |
@@
| 2023 | CAMEL (Li et al., 2023b) | None | Pre-S | Director |
Agent |
@@
| 2023 | Aher et al. (Aher et al., 2023) | R-P | Pre-S | Prototype |
Agent |
@@
| 2023 | CGMI (Jinxin et al., 2023) | S-P | Pre-S | Scriptwriter |
Env |
@@
| 2023 | ChatLaw (Cui et al., 2024b) | None | Pre-S | Director |
Agent |
@@
| 2020 | Alfred (Shridhar et al., 2020) | R-P | Pre-S | Director |
Agent |
@@
| 2023 | ChatDev (Qian et al., 2024) | None | Pre-S | Director |
Agent |
@@
| - | BactoWars (Berryman, 2008) | S-P | Pre-S | Scriptwriter |
Agent |
@@
| - | EINSTein (Berryman, 2008) | S-P | Pre-S | Scriptwriter |
Agent |
@@
| - | MANA (Berryman, 2008) | S-P | Pre-S | Scriptwriter |
Env |
@@
| 2005 | MASON (Luke et al., 2005) | S-P | Pre-S | Scriptwriter |
Agent |
@@
| 1999 | NetLogo (Wilensky, 1999) | S-P | Pre-S | Scriptwriter |
Agent |
@@
| 2006 | North et al. (North et al., 2006) | S-P | Pre-S | Scriptwriter |
Agent |
@@
| 1996 | Minar et al. (Minar et al., 1996) | S-P | Pre-S | Scriptwriter |
Agent |
@@
| 2023 | ComplexWorld (Basavatia et al., 2023) | S-V | Pre-S | Scriptwriter |
Env |
@@
| 2013 | Gaube et al. (Gaube and Remesch, 2013) | R-P | Pre-S | Prototype |
Agent |
@@
| 2023 | WarAgent (Hua et al., 2024) | S-P | Pre-S | Scriptwriter |
Agent |
@@
| 2018 | Kavak et al. (Kavak et al., 2018) | S-P | Pre-S | Prototype |
Agent |
@@
| 2023 | Modelscope-agent (Li et al., 2023a) | R-V | Pre-S | Scriptwriter |
Agent |
@@
| 2024 | AgentCoord (Pan et al., 2024) | S-P | Pre-S | Director |
Agent |
@@
| 2023 | Choicemates (Park et al., 2023a) | None | Pre-S | Director |
Agent |
@@
| 2024 | Rehearsal (Shaikh et al., 2024) | None | Pre-S | Director |
Env |
@@
| 2023 | Wang et al. (Wang et al., 2024c) | S-V | Pre-S | Scriptwriter |
Agent |
@@
| 2023 | Humanoid Agents (Wang et al., 2023a) | S-P | Pre-S | Scriptwriter |
Agent |
@@
| - | PedSim (PedSim, [n. d.]) | S-P | Pre-S | Scriptwriter |
Agent |
@@
| - | AnyLogic (Borshchev, 2014) | S-P | Pre-S | Scriptwriter |
Agent |
@@
| 2023 | AutoGPT (Significant Gravitas, [n. d.]) | None | Pre-S | Scriptwriter |
Agent |
@@
| 2023 | BabyAGI (BabyAGI, [n. d.]) | None | Pre-S | Scriptwriter |
Agent |
@@
| 2023 | CommunityBots (Jiang et al., 2023) | None | Pre-S | Director |
Agent |
@@
| 2024 | ComPeer (Liu et al., 2024b) | None | Pre-S | Director |
Agent |
@@
| 2024 | PrISM-Observer (Arakawa et al., 2024) | R-P | Pre-S | Director |
Agent |
@@
| 2024 | Jaber et al. (Jaber et al., 2024) | R-P | Pre-S | Director |
Agent |
@@
| 2024 | Chen et al. (Chen et al., 2024) | None | Pre-S | Director |
Agent |
@@
| 2024 | Zhang et al. (Zhang et al., 2024d) | R-V | Pre-S | Director |
Agent |
@@
| 2024 | Text2AC (Sun et al., 2024) | R-V | Pre-S | Director |
Agent |
@@
| 2023 | Ross et al. (Ross et al., 2023) | None | Pre-S | Director |
Agent |
@@
| 2024 | ChatCam (Xiao et al., 2024) | R-P | Pre-S | Director |
Agent |
@@
| 2024 | DrHouse (Yang et al., 2024) | R-P | Pre-S | Director |
Agent |
@@
| 2024 | ChatIoT (Gao et al., 2024b) | R-P | Pre-S | Director |
Agent |
@@
| 2024 | CrowdBot (Wang et al., 2024b) | R-P | Pre-S | Director |
Agent |
@@
| 2022 | Social Simulacra (Park et al., 2022) | S-V | D-S | Scriptwriter | Agent |
@@
| 2012 | Prom Week (McCoy et al., 2012) | S-P | D-S | Director | Agent |
@@
| 2011 | Prom Week (McCoy et al., 2011) | S-P | D-S | Director | Agent |
@@
| 2023 | AGENTVERSE (Chen et al., 2023) | S-P | D-S | Director | Sim |
@@
| 2023 | ChatEval (Chan et al., 2023) | None | D-S | Director | Agent |
@@
| 2024 | Zhang et al. (Zhang et al., 2024c) | S-P | D-S | Actor | Agent |
@@
| 2023 | Memory sandbox (Huang et al., 2023) | None | D-S | Scriptwriter | Agent |
@@
| 2022 | Inner monologue (Huang et al., 2022) | S-P; R-P | D-S | Director | Agent |
@@
| 2022 | Krishna et al. (Krishna et al., 2022) | R-V | Pre-S | Director | Agent |
@@
| 2023 | RoCo (Mandi et al., 2023) | R-P | D-S | Actor | Agent |
@@
| 2023 | ChatLaw (Cui et al., 2024b) | None | D-S | Director | Agent |
@@
| 2023 | Mehta et al. (Mehta et al., 2024) | S-P | D-S | Director | Agent |
@@
| 2020 | Alfred (Shridhar et al., 2020) | R-P | D-S | Director | Agent |
@@
| 2023 | Mohanty et al. (Mohanty et al., 2023) | S-P | D-S | Director | Agent |
@@
| 2022 | Teach (Padmakumar et al., 2022) | R-P | D-S | Director | Agent |
@@
| 2005 | MASON (Luke et al., 2005) | S-P | D-S | Director | Sim |
@@
| 2006 | North et al. (North et al., 2006) | S-P | D-S | Director | Sim |
@@
| 2023 | Drive like a human (Fu et al., 2023) | R-P | D-S | Director | Agent |
@@
| 2006 | Guyot et al. (Guyot and Honiden, 2006) | R-V | D-S | Actor | Agent |
@@
| 2023 | Surrealdriver (Jin et al., 2024b) | S-P | Pre-S | Director | Agent |
@@
| 2023 | The SocialAI School (Kovač et al., 2023) | S-P | Pre-S | Director | Agent |
@@
| 2023 | Choicemates (Park et al., 2023a) | None | D-S | Director | Agent |
@@
| 2023 | Ghost in the minecraft (Zhu et al., 2023) | R-V | D-S | Director | Agent |
@@
| 2024 | Lu et al. (Lu et al., 2024a) | None | D-S | Director | Agent |
@@
| 2024 | Teach AI How to Code (Jin et al., 2024a) | None | D-S | Director | Agent |
@@
| 2024 | Zhou et al. (Zhou et al., 2024a) | None | D-S | Director | Agent |
@@
| 2022 | Wordcraft (Yuan et al., 2022) | None | D-S | Director | Agent |
@@
| 2023 | Ross et al. (Ross et al., 2023) | None | D-S | Director | Sim |
@@
| 2023 | Generative Agents (Park et al., 2023b) | S-P | Post-S | Actor | Agent |
@@
| - | - | - | Post-S | Observer | Agent |
@@
| 2022 | Social Simulacra (Park et al., 2022) | S-V | Pre-S | Scriptwriter | Agent |
@@
| - | - | - | Post-S | Observer | Agent |
@@
| 2024 | Zhang et al. (Zhang et al., 2024c) | S-P | D-S | Actor | Agent |
@@
| 2023 | MetaGPT (Hong et al., 2024) | None | Post-S | Observer | Agent |
@@
| 2023 | Argyle et al. (Argyle et al., 2023) | R-P | Post-S | Observer | Agent |
@@
| 2023 | AgentSims (Lin et al., 2023) | S-P | D-S | Observer | Agent |
@@
| 2023 | Saha et al. (Saha et al., 2023) | None | D-S | Actor | Agent |
@@
| 2023 | CAMEL (Li et al., 2023b) | None | Post-S | Observer | Agent |
@@
| - | BactoWars (Berryman, 2008) | S-P | Post-S | Observer | Agent |
@@
| 2022 | FAIR et al. († et al.(2022), FAIR) | R-V | D-S | Actor | Agent |
@@
| 2020 | Feng et al. (Feng et al., 2020) | R-P | Pre-S | Prototype | Agent |
@@
| 2023 | Hämäläinen et al. (Hämäläinen et al., 2023) | R-P | Post-S | Observer | Agent |
@@
| 2023 | War and Peace (Hua et al., 2024) | S-P | Post-S | Observer | Agent |
@@
| 2023 | Surrealdriver (Jin et al., 2024b) | S-P | Post-S | Observer | Agent |
@@
| 2023 | Modelscope-agent (Li et al., 2023a) | R-V | Pre-S | Observer | Agent |
@@
| 2023 | Liu et al. (Liu et al., 2023) | S-P | Post-S | Observer | Agent |
@@
| 2023 | Alympics (Mao et al., 2024) | S-V | Post-S | Observer | Agent |
@@
| 2024 | AgentCoord (Pan et al., 2024) | S-P | D-S | Director | Agent |
@@
| 2023 | Choicemates (Park et al., 2023a) | None | D-S | Observer | Agent |
@@
| - | - | - | Post-S | Observer | Agent |
@@
| 2024 | Schwitzgebel et al. (Schwitzgebel et al., 2024) | None | Post-S | Observer | Agent |
@@
| 2024 | Rehearsal (Shaikh et al., 2024) | None | D-S | Director | Agent |
@@
| 2023 | Wang et al. (Wang et al., 2024c) | S-V | Post-S | Observer | Agent |
@@
| 2023 | Humanoid Agents (Wang et al., 2023a) | S-P | D-S | Observer | Agent |
@@
| - | - | - | Post-S | Observer | Agent |